### v1

# PoseBench3D: A Cross-Dataset Analysis Framework for 3D Human Pose Estimation

Abstract
Reliable three-dimensional human pose estimation is becoming increasingly important for real-world applications, yet much of prior work has focused solely on the performance within a single dataset. In practice, however, systems must adapt to diverse viewpoints, environments, and camera setups—conditions that differ significantly from those encountered during training, which is often the case in real-world scenarios. To address these challenges, we present a standardized testing environment in which each method is evaluated on a variety of datasets, ensuring consistent and fair cross-dataset comparisons—allowing for the analysis of methods on previously unseen data. Therefore, we propose PoseBench3D, a unified framework designed to systematically re-evaluate prior and future models across four of the most widely used datasets for human pose estimation—with the framework able to support novel and future datasets as the field progresses. Through a unified interface, our framework provides datasets in a pre-configured yet easily modifiable format, ensuring compatibility with diverse model architectures. We re-evaluated the work of 18 methods, either trained or gathered from existing literature, and reported results using both Mean Per Joint Position Error (MPJPE) and Procrustes Aligned Mean Per Joint Position Error (PA-MPJPE) metrics, yielding more than 100 novel cross-dataset evaluation results. Additionally, we analyze performance differences resulting from various pre-processing techniques and dataset preparation parameters—offering further insight into model generalization capabilities.
1 Introduction
Three-dimensional Human Pose Estimation (HPE) has gathered substantial interest for it’s critical role in applications such as healthcare [36, 29], action recognition [17, 16], military operations [46, 12], human-computer interaction [30, 38], and virtual/augmented reality [23]—among many others. Despite remarkable progress in recent years [14, 31, 47, 28, 9, 33], most work focuses on performance within a single, controlled dataset. The majority of existing work adopts either a direct-from-image approach or a two-stage approach, with the latter being widely used for its flexibility. In the two-stage approach, the first stage detects 2D keypoints, and the second stage, known as the “lifting network” [22, 42, 15, 4, 49, 11, 40] maps the 2D keypoints to 3D pose estimations. However, these two-stage models show poor generalization across various different datasets [35, 20].
(a) Elevation
(b) Azimuth
Figure 1: Elevation and azimuth distribution across four datasets – H36M, GPA, 3DPW, SURREAL. Elevation and azimuth are calculated relative to the subject’s local axis (a) The elevation distribution. Notice how all the datasets have similar elevation profiles. (b) The azimuth distribution for the same configuration. We see stark differences across all datasets.
To assess the generalization of these models, we consider cross-dataset performance: the task of training on one dataset and testing on another. Machine learning models are generally susceptible to overfitting to the bias and variance of the data distribution it was trained on; this is particularly pronounced in 3D human pose estimation datasets due to variations in data collection procedures, where the underlying distributions of datasets often differ significantly. For example, the H36M dataset [10] was captured using a four-camera setup, while GPA [34] used five, and both 3DPW [21] and SURREAL [32] relied on a single-camera configuration. In fig. 1, we show the frequency histogram of camera positions relative to the subject in spherical coordinates. While elevation distributions are largely consistent across datasets (cf. fig. 1(a)), the azimuthal angles vary significantly (cf. fig. 1(b)).
Beyond camera setup, joint placement conventions also differ between datasets. For instance, there is currently no standardized convention for placing the hip or spine joints (cf. fig. 2)—note the difference in width of hips in fig. 2(a) and fig. 2(c). Additionally, since datasets are collected with varying action sets, the captured range of motion also varies. In fig. 3, the blue dots represent the recorded range of motion, whereas the yellow region shows the valid range. Since these two datasets do not explore the entire valid range of motion, normalization performed with one will not translate well onto the other.
Research Gap. These vast differences mean that, although any method may achieve near-optimal performance within its own dataset, it is likely to perform poorly across differing datasets. Measuring cross-dataset performance is paramount, but is extremely labor-intensive to perform manually. Our goal in this paper is to automate the task of cross-dataset comparison.
(a) H36M
(b) GPA
(c) 3DPW
(d) SURREAL
Figure 2: Sample pose data points from H36M, GPA, 3DPW, and SURREAL. Compared to 3DPW, H36M shows a wider hip joint spread and more upright head posture.
(a) H36M
(b) 3DPW
Figure 3: Scatter plots showing limb range of motion in H36M and 3DPW datasets. Blue represents regions of motion observed in the datasets, yellow represents the full range of motion based on physical limitations of limbs, and purple represents regions that cannot be reached without physical injury to the limb. Note the distribution differences between datasets.
Proposed Work. To advance the field and encourage broader generalization, methods and architectures must be evaluated on multiple datasets under a consistent evaluation protocol—in order to better facilitate and automate cross-dataset evaluation. In this spirit, we introduce a unified test bench that consolidates several 3D human pose datasets within a standardized testing environment. This approach enables detailed analysis of generalization capabilities, method scalability, and performance trends across diverse environmental conditions—mirroring realistic scenarios in which settings are rarely uniform. Our aim is to narrow the existing research gap by offering insights into pose estimation that are simply not possible when relying solely on single-dataset evaluations.
Our contributions are summarized as follows:
  1. 1.
    Standardized Benchmarking Environment. We introduce a standardized evaluation framework, PoseBench3D, that consolidates four commonly used human pose estimation datasets—H36M[10], GPA[34], 3DPW[21], and SURREAL[32]—while providing modularity for future datasets. Our framework simplifies cross-dataset evaluation by requiring, simply, a single user-supplied configuration file (cf. section A.1) specifying model, experiment, and user customization details.
  2. 2.
    Re-Evaluation of 18 Methods. We release detailed cross-dataset evaluation metrics for 18 established methods, generating over 100 previously unreported comparisons—shedding light on the need for standardized, comprehensive benchmarking in 3D human pose estimation.
  3. 3.
    Open and Extensible Design. PoseBench3D is structured to accommodate both new models and datasets as the field progresses, encouraging fair, reproducible, and easily customizable evaluations through our open-source framework. We make our code publicly available.
  4. 4.
    Detailed Analysis. We also investigate the impact of factors such as viewpoint distribution and z-score standardization on a model’s ability to generalize across datasets.
2 Related Work
Current Cross-Dataset Evaluation. While several methods have attempted to address generalization, none have demonstrated success beyond testing on a single additional dataset beyond their training set. In particular, TCPFormer [14], FinePOSE [37], MotionBERT [48], and MotionAGFormer [28] all train on Human3.6M [10] and test only on 3DPW [21]. Although each introduces novelty—TCPFormer in lifting, FinePOSE in architectural design, and MotionBERT and MotionAGFormer in binary decisions regarding the nature of the test image—none have demonstrated the ability to generalize across multiple datasets, let alone real-world scenarios. In contrast, PoseBench3D automates cross-dataset evaluation, eliminating long-existing friction in benchmarking across datasets and addressing a key limitation that has held back progress in the field.
Shortage of Testing Environments. Although we commend AdaptPose [7] for evaluating its fine-tuned version of [24] and [8] across four datasets—Human3.6M, 3DHP, 3DPW, and Ski-Pose—they do not provide a standardized evaluation framework for others to test their own models. Wang et al. [35], Manzur et al. [20], and Gong et al. [8] also report cross-dataset results, which aligns more closely with our goals. Nevertheless, there remains a lack of standardized cross-dataset evaluation frameworks that enable consistent benchmarking across multiple works.
3 Overview of PoseBench3D
Our framework, illustrated in the high-level overview in Figure 4, is designed to support the evaluation of 2D-to-3D human pose lifting networks. In what follows, we provide a brief summary of the supported 2D-to-3D pose estimation datasets, outline the curated set of benchmark models included in our evaluation, and describe the techniques employed to standardize these datasets under a unified and fair comparison framework.
Figure 4: High-level overview of interactions among the framework components.
3.1 Benchmark Datasets
Our framework currently supports four datasets for evaluation, depicted in table 1. The Human3.6M dataset (H36M) [10] consists of over 3.6 million annotated 3D human poses from 15 natural activity scenarios. This dataset is currently the most popular for use in the field of 3D human pose estimation, but its controlled lab environment limits its generalization potential. Geometric Pose Affordance (GPA) [34] focuses on the interaction between humans and objects. With 305 thousand RGB images, GPA emphasizes environmental interaction. The 3D Poses in the Wild (3DPW) dataset [21] emphasizes real-world, unconstrained environments with pose data captured using IMU sensors along with automatically generated 2D-to-3D pose associations. SURREAL [32] is a large synthetic dataset for 3D pose estimation, derived from the CMU Motion Capture database [5]. It uses MoSh [19] to match SMPL [18] parameters to raw marker paths. Since the dataset is synthetic and uses human skeletal models, both 2D and 3D ground truth joint locations are available to arbitrary precision.
Table 1: Comparative summary of four benchmark 3D human pose datasets, highlighting differences in camera systems, data modalities, and geometric attributes. Each dataset is color coded through out the paper for ease of comparison.
Datasets
Criteria Human3.6M GPA 3DPW SURREAL
No. of Cameras 4 RGB + 1 TOF 2 RGB + 3 RGBD 1 Moving Camera 1 Virtual Camera
Subjects & Activities 11 Actors, 15 Daily Actions 13 Subjects, Scripted Interactions 7 Subjects, 18 Clothing Types Synthetic, 6.5 M Frames from CMU MoCap
Environment Indoor Lab Indoor Studio, Rich 3D Scenes Outdoor Real-World Settings Synthetic Indoor Scenes
Imaging Space $1000\times 1002$ $1920\times 1080$ $1920\times 1080$ $320\times 240$
Camera Distance (m) $5.2\pm 0.8$ $5.1\pm 1.2$ $3.5\pm 0.7$ $8.0\pm 1.0$
Camera Height (m) $1.6\pm 0.05$ $1.0\pm 0.3$ $0.6\pm 0.8$ $0.9\pm 0.1$
Focal Length (mm) $1146.8\pm 2.0$ $1172.4\pm 121.3$ $1962.2\pm 1.5$ $600\pm 0$
No. of Joints 38 34 24 24
3.2 Supported Model Architectures
Our framework is set-up to support Pose Lifting Networks models out-of-the-box, with the goal of supporting video- and image-based models in future work. 2D-to-3D pose estimation models, a.k.a. “Lifting Networks”, are widely used for their flexibility, augmentation capabilities, and ease of implementation. Given that 2D keypoint-based representations lack many visual cues present in images (such as occlusion), it simplifies the task of 3D pose estimation–often at the cost of generalization. As reported by [35, 20], they show poor performance when tested against other datasets. Therefore, we survey the lifting networks reported by previous works [22, 42, 44, 39, 41, 1, 6, 8, 45, 43, 40, 4, 49] and pick 18 configurations based on their availability and adaptability. Gong et al. [8] reported results on Martinez et al. [22], Zhao et al. [42], Pavllo et al. [25], and Cai et al. [1] with optimized and un-optimized variants tested on the H36M and 3DPW datasets. Since the checkpoints are already provided, we used these checkpoints to re-evaluate the models on all four datasets with our framework. We also gathered five transformer-based models [45, 43, 44, 26, 13], two graph convolution models [39, 42], a spatio-temporal encoder model [41], and a diffusion-based model [27]. Given that various models from the authors work were missing checkpoints or were trained on 17 joint skeleton configurations, we decided to standardize testing by retraining these models from scratch on the more common 16 joint skeleton configuration to remain fair and consistent across comparisons. We further ensure the retrained model is as good as the original one by comparing the reported same-dataset results directly.
3.3 Dataset Preparation
To unify four distinct datasets under a common processing interface, we first identified the core elements required by lifting-based 3D pose estimation networks. These models typically take 2D keypoints as input and predict 3D joint coordinates as output. The 3D pose is usually expressed either in world-space or camera-space, with the latter being both statistically justified and widely adopted due to its alignment with image-space supervision. Additionally, most models standardize the pose by centering the skeleton at the hip joint. As shown in table 1 and discussed in [35], datasets differ significantly in terms of joint definitions, camera intrinsics/extrinsics, image resolution, and other parameters. As such, each dataset requires tailored preprocessing pipelines. For efficiency, we preprocessed and cached all dataset files as NumPy zipped archives, enabling faster data loading. Although different datasets call for different preprocessing techniques, the most important operation is acquiring and converting 3D joint positions into camera coordinates. Once this is established, projecting to 2D space follows a standardized procedure.
The H36M [10] dataset provides four cameras with their intrinsic and extrinsic matrices. While the intrinsic matrices stay the same for all 15 action sequences across 11 subjects, the extrinsic parameters (e.g. orientation and translation) change. There are 15 actions, 2 subactions per action, and 4 cameras, all acted by 11 subjects. Since the 3D pose estimation task focuses training on subjects 1, 5, 6, 7, and 8 and testing on subjects 9 and 11, we only provide data points for these subjects. To avoid redundancy, we only sample frames that have noticeable change in the pose. The world-space 3D coordinates ($\mathrm{X}\in\mathbb{R}^{N\times 3}$) are then converted into camera-space coordinates with the help of $(\mathrm{R}\cdot(\mathrm{X}^{\top}-\mathrm{t}))^{\top}$, where $\mathrm{R}\in\mathbb{R}^{3\times 3}$ is the rotation matrix and $\mathrm{t}$ is the translation vector. The 2D points were projected using the perspective projection equation to obtain the ground-truth 2D keypoints. On the other hand, the GPA [34] dataset provides intrinsic and extrinsic parameters for every frame where all the annotations are provided in a json file. The key difference in processing the dataset was the camera-space coordinate transformation. We use $\mathrm{R_{i}}^{\top}\cdot\mathrm{X}_{i}^{\top}+\mathrm{t_{i}}$, where the notification follows a similar convention to that of the H36M dataset. The translation vector provided was in centimeters in the original dataset, where the 3D coordinates were found to be in millimeters. The rotation matrix was also stored in a vector format. The 3DPW dataset [21] immediately differentiates itself from the rest of the datasets in terms of formatting. The dataset provides camera extrinsics and intrinsics per sequence, and each sequence can contain multiple subjects. We process each subject separately. Since the camera parameters are provided as an extrinsic matrix ($\mathbb{R}^{4\times 4}$), a multiplication with the world space coordinates expressed as homogenous coordinates is enough to obtain the camera space coordinates. The SURREAL dataset [32] required minimal preprocessing for camera coordinate conversion. However, we identified several invalid data batches, which consistently caused large pose estimation errors. To maintain consistency in evaluation, we excluded these anomalous samples, which accounted for less than 1% of the dataset.
3.4 Framework Initialization
[…]
4 Experiments and Results
4.1 Setup
Not all datasets agree on a common seventeenth joint. Therefore, we retrain models originally trained with a 17-joint configuration using only 16 joints. Following established conventions in the field [2, 22, 42, 3, 4], we report both Protocol 1 error—MPJPE (Mean Per-Joint Position Error)—and Protocol 2 error—PA-MPJPE (MPJPE after Procrustes Alignment), commonly referred to as Protocol 1 and Protocol 2, respectively. Since our experiments focus on lifting networks, they are not memory intensive. All experiments were conducted using two A30 24GB GPUs. Note that our framework supports both GPU and CPU systems.
4.2 Cross-Dataset Evaluation
We report the MPJPE scores in table 2 and PA-MPJPE scores in table 3. We also include two cross-dataset results available from prior work marked with $\ddagger$ in the tables.
[…]
Table 2: Cross-dataset evaluation sorted by decreasing average MPJPE (mm). All models are trained on H36M. Lower is better. $\dagger$: PoseAug-optimized; $\diamond$: unoptimized variant; $\bullet$: retrained from scratch; $\ddagger$: reported from prior work.
Cross-Dataset Evaluation
Model Name H36M GPA 3DPW SURREAL Average Error (MPJPE $\downarrow$)
GraFormer [44] $\bullet$ 36.44 259.11 308.96 150.46 188.74
SEM-GCN [8] $\diamond$ 47.03 262.34 315.31 118.30 185.75
SEM-GCN [8] $\dagger$ 41.90 241.21 239.07 107.26 157.36
VideoPose [8] $\diamond$ 41.47 208.45 257.81 107.96 153.92
GLA-GCN [39] $\bullet$ 44.51 237.29 207.59 119.08 152.12
ST-GCN [8] $\diamond$ 41.52 205.76 238.48 107.61 148.34
Martinez et al. [8] $\diamond$ 41.42 205.62 226.20 110.01 145.81
PoseFormer V1 [45] $\bullet$ 42.82 217.90 161.97 156.59 144.82
PoseFormer V2 [43] $\bullet$ 42.80 209.90 162.45 146.39 140.39
KTPFormer [26] $\bullet$ 38.12 205.71 193.63 108.95 136.60
MixSTE [41] $\bullet$ 38.44 182.13 171.28 131.23 130.77
D3DP [27] $\bullet$ 39.61 189.74 148.56 127.90 126.45
ST-GCN [8] $\dagger$ 36.83 185.63 174.16 101.95 124.64
MHFormer [13] $\bullet$ 42.60 202.59 202.59 114.55 120.86
Martinez et al. [8] $\dagger$ 39.11 169.79 134.12 98.99 110.50
VideoPose [8] $\dagger$ 39.02 174.39 126.05 100.42 109.97
Wang et al. [35] $\ddagger$ 52.00 98.30 109.5 114.00 93.45
Manzur et al. [20] $\ddagger$ 33.52 92.31 95.83 65.62 71.82
[…]
Table 3: Cross-dataset evaluation sorted by decreasing average PA-MPJPE (mm). All models are trained on H36M. Lower is better. $\dagger$: PoseAug-optimized; $\diamond$: unoptimized variant; $\bullet$: retrained from scratch; $\ddagger$: reported from prior work.
Cross-Dataset Evaluation
Model Name H36M GPA 3DPW SURREAL Average Error (PA-MPJPE $\downarrow$)
SEM-GCN [8] $\diamond$ 36.12 178.43 166.76 87.35 117.17
GraFormer [44] $\bullet$ 28.40 152.31 189.30 87.87 114.47
ST-GCN [8] $\diamond$ 32.47 125.99 206.45 69.11 108.51
SEM-GCN [8] $\dagger$ 33.66 166.88 131.38 80.98 103.23
PoseFormer V1 [45] $\bullet$ 33.50 138.12 103.89 95.72 92.81
GLA-GCN [39] $\bullet$ 35.27 148.26 106.10 73.31 90.74
PoseFormer V2 [43] $\bullet$ 33.18 145.94 92.47 91.02 90.65
KTPFormer [26] $\bullet$ 30.27 133.73 127.09 67.66 89.69
ST-GCN [8] $\dagger$ 28.69 112.12 131.99 65.17 84.49
Martinez [8] $\diamond$ 31.80 124.17 111.30 67.97 83.81
VideoPose [8] $\diamond$ 32.17 126.61 102.65 65.67 81.78
MixSTE [41] $\bullet$ 31.05 120.85 91.12 76.96 80.00
D3DP [27] $\bullet$ 31.25 133.82 73.86 78.83 79.44
MHFormer [13] $\bullet$ 32.56 124.50 124.50 69.18 73.21
Martinez [8] $\dagger$ 30.31 103.26 79.74 59.94 68.31
VideoPose [8] $\dagger$ 30.17 108.92 75.29 58.33 68.18
Manzur et al. [20] $\ddagger$ 29.10 69.48 64.28 51.53 53.60
Figure 5: Per-joint errors for Protocols 1 and 2, averaged across all models and datasets.
We also analyze per-joint position errors under both protocols in fig. 5, averaged across all 18 model configurations and all four datasets. While the endpoints of the human body—such as the wrists, ankles, and head—are more susceptible to errors, note that Procrustes alignment alone results in a substantial reduction in these errors (highlighted in red). Since Procrustes alignment accounts for rigid transformations, this highlights that human “pose” estimation is distinct from “position” estimation. Therefore, future efforts should focus on reducing angular errors in the bones rather than solely minimizing positional discrepancies.
4.3 Impact of Z-Score Standardization
Z-score standardization is a common operation to normalize the data by zero-centering with the mean and scaling by the standard deviation. In the context of pose estimation, where datasets may originate from diverse distributions, such normalization indeed improves cross-dataset performance when the mean and standard deviation are known—though they are typically unavailable during inference time in real-world scenarios. Table 4 shows the models [22] and [42] trained with H36M and normalized with the mean and standard deviation of the same dataset. For the other datasets, we used their respective mean and standard deviation to normalize them. We also include the unoptimized ($\diamond$) and PoseAug-optimized ($\dagger$) variants from table 2 for comparison, and also note the percentage improvement relative to the unoptimized variant in both cases. From all the cases, note that z-score normalization performed with the test set’s mean and standard deviation improves cross-dataset generalization significantly for both models (e.g., SemGCN sees an improvement of more than 50% in 3DPW). PoseAug’s optimization adds novel poses to the training and secures a place right in the middle of the unoptimized variant and the z-score standardized variant. This shows that models overfit on the distribution they are trained on—which can often be countered by increasing variation.
Table 4: MPJPE comparison of Martinez and SEM-GCN models across datasets. Percentage improvements relative to the baseline ($\diamond$) are in parentheses.
(a) Martinez Comparison
Model H36M GPA 3DPW SURREAL Average Error
Unoptimised $\diamond$ 41.42 205.62 226.20 110.01 145.81
Optimised $\dagger$ 39.11 ($\downarrow$ 5.6%) 169.79 ($\downarrow$ 17.4%) 134.12 ($\downarrow$ 40.7%) 98.99 ($\downarrow$ 10.0%) 110.50 ($\downarrow$ 24.2%)
Z-score 52.37 ($\uparrow$ 26.4%) 104.39 ($\downarrow$ 49.3%) 141.10 ($\downarrow$ 37.6%) 81.64 ($\downarrow$ 25.8%) 94.88 ($\downarrow$ 34.9%)
(b) SEM-GCN Comparison
Model H36M GPA 3DPW SURREAL Average Error
Unoptimised $\diamond$ 47.03 262.34 315.31 118.30 185.75
Optimised $\dagger$ 41.90 ($\downarrow$ 10.9%) 241.21 ($\downarrow$ 8.1%) 239.07 ($\downarrow$ 24.2%) 107.26 ($\downarrow$ 9.3%) 157.36 ($\downarrow$ 15.3%)
Z-score 53.94 ($\uparrow$ 14.7%) 114.85 ($\downarrow$ 56.2%) 153.61 ($\downarrow$ 51.3%) 99.88 ($\downarrow$ 15.6%) 105.57 ($\downarrow$ 43.2%)
4.4 Impact of Viewpoint
One interesting experiment is to look at the interplay between the viewpoint distribution of the datasets and its impact on the MPJPE error. In figs. 6 and 7, we present the MPJPE scores against the viewpoint distribution over the training and testing sets simultaneously. One common trend across all these figures is that whenever the viewpoint distribution disagrees with the training set’s viewpoint distribution, we observe a spike in error. This shows how much critical viewpoint is in 3D human pose estimation. Table 5 formally measures the inverse relationship between training data frequency and joint error by computing the Spearman correlation between the two, measured across patches on the viewpoint sphere measuring 5°5  ° $\times$ 10°10°10  ° degrees in elevation and azimuth, respectively. The correlation is always negative with high statistical significance. Notice how the p-value is lower when the evaluation is cross-dataset, indicating the variance in viewpoint distribution severely impacts performance. Please refer to Section A.2 for more figures.
(a) H36M
(b) GPA
(c) 3DPW
(d) SURREAL
Figure 6: MPJPE score (SEM-GCN [42] trained on H36M) compared against viewpoint distribution in azimuth relative to the subject.
(a) H36M
(b) GPA
(c) 3DPW
(d) SURREAL
Figure 7: MPJPE score (SEM-GCN [42] trained on H36M) compared against viewpoint distribution in elevation relative to the subject.
Table 5: Spearman correlation (which contained at least 5 training images and at least 5 test images) between training viewpoint distributions and test error (5°5  ° elev × 10°10°10  ° azim bins). Num = number of bins; sigma is the p-value represented as the number of standard deviations from random.
Num Train Test Spearman p-val sigma
[…]
380 SURREAL H36M -0.19 0.018 380.018380.018\,380.018 38 3.70
[…]
417 H36M 3DPW -0.20 0.003 7010.0037010.003\,7010.003 701 4.13
[…]
5 Conclusion and Future Work
We introduced a framework for evaluating cross-dataset performance of 2D-to-3D pose lifting networks. PoseBench3D enables standardized, reproducible evaluation and supports automatic benchmarking across multiple datasets. We apply PoseBench3D to evaluate 18 configurations of models, using checkpoints either obtained from prior work or retrained according to the original authors’ specifications. With more than 100 newly reported cross-dataset comparisons, we analyzed the results through several lenses and identify key factors—such as viewpoint distribution and normalization strategies—that significantly impact generalization in 3D human pose estimation.
The limitation of this work is that the current framework supports only four widely used datasets and is limited to pose lifting networks. In the future, we plan to extend PoseBench3D to support image-based models, enabling evaluation of end-to-end systems that go from RGB input to 3D pose. We also aim to integrate additional datasets and enhance the modularity of the framework to further encourage community contributions and broader adoption. Ultimately, we hope PoseBench3D will serve as a foundation for fair, comprehensive, and scalable benchmarking in 3D human pose estimation.
Appendix A Experiments
A.1 Example User-Supplied Config File
Listing 1: User-Supplied YAML Config File
# Unified YAML configuration
# General settings
dataset: [h36m, surreal, 3dpw, gpa]
keypoints: ["gt": ground truth]
checkpoint: "checkpoint"
evaluate: "/path_to_checkpoint.pt"
model_type: [JIT-Traced PyTorch File, ONNX TensorFlow File]
# Dataset locations
path_to_h36m: path_to_h36m
path_to_gpa: path_to_gpa
path_to_surreal_train: path_to_surreal_train
path_to_surreal_test: path_to_surreal_test
path_to_3dpw: path_to_3dpw
data_dir: path_to_data_directory
# Runtime & miscellaneous options
num_workers: 2-4
print_sample: [True, False]
per_joint_error: [True, False]
save_predictions: [True, False]
# Input tensor shape
input_shape:
batch_size: [1…N]
num_frames: 1
num_joints: [14, 16]
coordinates: 2
# Output tensor shape
output_shape:
batch_size: [1…N]
num_frames: 1
num_joints: [14, 16]
coordinates: 3
# Model-specific flags
model_info:
flattened_coordinate_model: [True, False]
output_3d: [Meters, Millimeters]
root_center_2d_test_input: [True, False]
normalize_2d_to_minus_one_to_one: [True, False]
trained_on_normalized_data: [True, False]
video_model: [True, False]
A.2 Cross-Dataset Evaluation Visualization
Figure 8: Protocol 1 visual results from table 2 shown in increasing average dataset error.
Figure 9: Protocol 2 visual results from table 3 shown in increasing average dataset error.
Figure 10: Per-joint errors for Protocols 1 and 2, averaged across all models on H36M.
Figure 11: Per-joint errors for Protocols 1 and 2, averaged across all models on GPA.
Figure 12: Per-joint errors for Protocols 1 and 2, averaged across all models on 3DPW.
Figure 13: Per-joint errors for Protocols 1 and 2, averaged across all models on SURREAL.
(a) H36M
(b) GPA
(c) 3DPW
(d) SURREAL
Figure 14: MPJPE score (SEM-GCN [42] trained on GPA) compared against viewpoint distribution in azimuth relative to the subject.
(a) H36M
(b) GPA
(c) 3DPW
(d) SURREAL
Figure 15: MPJPE score (SEM-GCN [42] trained on GPA) compared against viewpoint distribution in elevation relative to the subject.
(a) H36M
(b) GPA
(c) 3DPW
(d) SURREAL
Figure 16: MPJPE score (SEM-GCN [42] trained on 3DPW) compared against viewpoint distribution in azimuth relative to the subject.
(a) H36M
(b) GPA
(c) 3DPW
(d) SURREAL
Figure 17: MPJPE score (SEM-GCN [42] trained on 3DPW) compared against viewpoint distribution in elevation relative to the subject.
(a) H36M
(b) GPA
(c) 3DPW
(d) SURREAL
Figure 18: MPJPE score (SEM-GCN [42] trained on GPA) compared against viewpoint distribution in elevation relative to the subject.
(a) H36M
(b) GPA
(c) 3DPW
(d) SURREAL
Figure 19: MPJPE score (SEM-GCN [42] trained on SURREAL) compared against viewpoint distribution in azimuth relative to the subject.
(a) H36M
(b) GPA
(c) 3DPW
(d) SURREAL
Figure 20: MPJPE score (SEM-GCN [42] trained on SURREAL) compared against viewpoint distribution in elevation relative to the subject.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Both the abstract and introduction accurately reflect the paper’s scope. We discuss the current friction and research gap, how we set out to address these issues, introduce a benchmarking evaluation environment to partner our approach, and provide results according to our proposed approach.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitation and future work in the conclusion section (section 5).
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not contain theoretical results; instead, it focuses on the practical applications of human pose estimation.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The paper outlines in detail the requirements necessary to reproduce our results. We mention all three components: an extensive overview of required datasets section 3.1, the models and architectures supported by the framework section 3.2, and the user-supplied configuration file for customization section A.1. We make our codebase available to the public for extension and reproduction of our work.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Our framework is fully open-access, with a link to its repository provided. The repository includes detailed instructions for dataset setup, model import, and configuration file requirements needed to reproduce and extend our results.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide details about our experimental settings/details in the github repository linked with the paper. While we mention the specifics of our testing environment, our framework does not support training. For model training details and settings, please refer to the original work by the authors.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We present the statistical significance of the evaluation results against viewpoint distribution through Spearman correlation (c.f. table 5).
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We present the compute resources in section 4.1.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our paper clearly details the code and resources used in this project to ensure both the reproducibility of our results and ease of extension for future work. All raw datasets are publicly available at the original sources provided by their respective authors. Additionally, we provide the post-processed data used in our experiments.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We highlight the positive impact of advancing 3D Human Pose Estimation, particularly in rapidly growing fields as noted in the Introduction. We do not believe there are any significant negative societal impacts associated with improving 3D Human Pose Estimation.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We expect that 3D Human Pose Estimation benchmarking and the datasets have a low risk for misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We use four datasets, all of which are free for research purposes and credited with appropriate citations. The eighteen model configurations that we use for experiments or re-evaluation are licensed under one of the following: CC-BY-4.0, MIT, Apache License 2.0. Some of the models do not mention any citation. Nevertheless, we cited all the models used in our experiments.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide the original code repository, checkpoints, and data preparation script with proper documentation.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our work does not make use of crowdsourcing or human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our work does not make use of human subjects
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: Our work does not make use of LLMs in a way that is important to the core methods of this paper.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@
-PoseBench3D: A Cross-Dataset Analysis Framework for 3D Human Pose Estimation
+PoseBench3D: A Cross-Dataset Analysis Framework for 3D Human Pose Estimation via Pose Lifting Networks
 Abstract
-Reliable three-dimensional human pose estimation is becoming increasingly important for real-world applications, yet much of prior work has focused solely on the performance within a single dataset.
+Reliable three-dimensional human pose estimation (3D HPE) remains challenging due to the differences in viewpoints, environments, and camera conventions among datasets.
+As a result, methods that achieve near-optimal in-dataset accuracy often degrade on unseen datasets.
 In practice, however, systems must adapt to diverse viewpoints, environments, and camera setups—conditions that differ significantly from those encountered during training, which is often the case in real-world scenarios.
-To address these challenges, we present a standardized testing environment in which each method is evaluated on a variety of datasets, ensuring consistent and fair cross-dataset comparisons—allowing for the analysis of methods on previously unseen data.
-Therefore, we propose PoseBench3D, a unified framework designed to systematically re-evaluate prior and future models across four of the most widely used datasets for human pose estimation—with the framework able to support novel and future datasets as the field progresses.
-Through a unified interface, our framework provides datasets in a pre-configured yet easily modifiable format, ensuring compatibility with diverse model architectures.
-We re-evaluated the work of 18 methods, either trained or gathered from existing literature, and reported results using both Mean Per Joint Position Error (MPJPE) and Procrustes Aligned Mean Per Joint Position Error (PA-MPJPE) metrics, yielding more than 100 novel cross-dataset evaluation results.
-Additionally, we analyze performance differences resulting from various pre-processing techniques and dataset preparation parameters—offering further insight into model generalization capabilities.
+Measuring cross-dataset performance is a vital process, but extremely labor-intensive when done manually for human pose estimation.
+To address these challenges, we automate this evaluation using PoseBench3D, a standardized testing framework that enables consistent and fair cross-dataset comparisons on previously unseen data.
+PoseBench3D streamlines testing across four widely used 3D HPE datasets via a single, configurable interface.
+Using this framework, we re-evaluate 18 methods and report over 100 cross-dataset results under Protocol 1: MPJPE and Protocol 2: PA-MPJPE, revealing systematic generalization gaps and the impact of common preprocessing and dataset setup choices.
+The PoseBench3D code is found at: https://github.com/bryanjvela/PoseBench3D.
 1 Introduction
-Three-dimensional Human Pose Estimation (HPE) has gathered substantial interest for it’s critical role in applications such as healthcare [36, 29], action recognition [17, 16], military operations [46, 12], human-computer interaction [30, 38], and virtual/augmented reality [23]—among many others.
+Three-dimensional Human Pose Estimation (HPE) has gathered substantial interest for its critical role in applications such as healthcare [37, 29], action recognition [17, 18], military operations [47, 13], human-computer interaction [30, 39], and virtual/augmented reality [23]—among many others.
 Despite remarkable progress in recent years [15, 31, 48, 28, 10, 34], most work focuses on performance within a single, controlled dataset.
@@
 To assess the generalization of these models, we consider cross-dataset performance: the task of training on one dataset and testing on another.
-Machine learning models are generally susceptible to overfitting to the bias and variance of the data distribution it was trained on; this is particularly pronounced in 3D human pose estimation datasets due to variations in data collection procedures, where the underlying distributions of datasets often differ significantly.
+Machine learning models are generally susceptible to overfitting to the bias and variance of the data distribution they were trained on; this is particularly pronounced in 3D human pose estimation datasets due to variations in data collection procedures, where the underlying distributions of datasets often differ significantly.
 For example, the H36M dataset [11] was captured using a four-camera setup, while GPA [35] used five, and both 3DPW [33] and SURREAL [32] relied on a single-camera configuration.
-In fig.
-1, we show the frequency histogram of camera positions relative to the subject in spherical coordinates.
+In Fig.˜1, we show the frequency histogram of camera positions relative to the subject in spherical coordinates.
 While elevation distributions are largely consistent across datasets (cf.
-fig.
-1(a)), the azimuthal angles vary significantly (cf.
-fig.
-1(b)).
+Fig.˜1(a)), the azimuthal angles vary significantly (cf.
+Fig.˜1(b)).
 Beyond camera setup, joint placement conventions also differ between datasets.
 For instance, there is currently no standardized convention for placing the hip or spine joints (cf.
-fig.
-2)—note the difference in width of hips in fig.
-2(a) and fig.
-2(c).
+Fig.˜2)—note the difference in width of hips in Fig.˜2(a) and Fig.˜2(c).
 Additionally, since datasets are collected with varying action sets, the captured range of motion also varies.
-In fig.
-3, the blue dots represent the recorded range of motion, whereas the yellow region shows the valid range.
+In Fig.˜3, the blue dots represent the recorded range of motion, whereas the yellow region shows the valid range.
 Since these two datasets do not explore the entire valid range of motion, normalization performed with one will not translate well onto the other.
 Research Gap.
-These vast differences mean that, although any method may achieve near-optimal performance within its own dataset, it is likely to perform poorly across differing datasets.
+These differences mean that, although any method may achieve near-optimal performance within its own dataset, it is likely to perform poorly across differing datasets.
 Measuring cross-dataset performance is paramount, but is extremely labor-intensive to perform manually.
@@
 We introduce a standardized evaluation framework, PoseBench3D, that consolidates four commonly used human pose estimation datasets—H36M[11], GPA[35], 3DPW[33], and SURREAL[32]—while providing modularity for future datasets.
-Our framework simplifies cross-dataset evaluation by requiring, simply, a single user-supplied configuration file (cf.
-section A.1) specifying model, experiment, and user customization details.
+Our framework simplifies cross-dataset evaluation by requiring, simply, a single user-supplied configuration file specifying model, experiment, and user customization details.
   2. 2.
@@
 2 Related Work
-Current Cross-Dataset Evaluation.
+Current Cross‑Dataset Evaluation.
 While several methods have attempted to address generalization, none have demonstrated success beyond testing on a single additional dataset beyond their training set.
-In particular, TCPFormer [14], FinePOSE [37], MotionBERT [48], and MotionAGFormer [28] all train on Human3.6M [10] and test only on 3DPW [21].
+In particular, TCPFormer [15], FinePOSE [38], and MotionAGFormer [28] all train on Human3.6M [11] and test only on 3DPW [33].
 Although each introduces novelty—TCPFormer in lifting, FinePOSE in architectural design, and MotionBERT and MotionAGFormer in binary decisions regarding the nature of the test image—none have demonstrated the ability to generalize across multiple datasets, let alone real-world scenarios.
@@
 Shortage of Testing Environments.
-Although we commend AdaptPose [7] for evaluating its fine-tuned version of [24] and [8] across four datasets—Human3.6M, 3DHP, 3DPW, and Ski-Pose—they do not provide a standardized evaluation framework for others to test their own models.
+Although we commend AdaptPose [8] for evaluating its fine-tuned version of [24] and [9] across four datasets—Human3.6M, 3DHP, 3DPW, and Ski‑Pose—they do not provide a standardized evaluation framework for others to test their own models.
 Wang et al.
@@
 3 Overview of PoseBench3D
-Our framework, illustrated in the high-level overview in Figure 4, is designed to support the evaluation of 2D-to-3D human pose lifting networks.
+Our framework, illustrated in the high-level overview in Figure˜4, is designed to support the evaluation of 2D-to-3D human pose lifting networks.
 In what follows, we provide a brief summary of the supported 2D-to-3D pose estimation datasets, outline the curated set of benchmark models included in our evaluation, and describe the techniques employed to standardize these datasets under a unified and fair comparison framework.
@@
 3.1 Benchmark Datasets
-Our framework currently supports four datasets for evaluation, depicted in table 1.
+Our framework currently supports four datasets for evaluation, depicted in Tab.˜1.
 The Human3.6M dataset (H36M) [11] consists of over 3.6 million annotated 3D human poses from 15 natural activity scenarios.
@@
 Geometric Pose Affordance (GPA) [35] focuses on the interaction between humans and objects.
-With 305 thousand RGB images, GPA emphasizes environmental interaction.
+With $\approx$ 305,000 RGB images, GPA emphasizes environmental interaction.
 The 3D Poses in the Wild (3DPW) dataset [33] emphasizes real-world, unconstrained environments with pose data captured using IMU sensors along with automatically generated 2D-to-3D pose associations.
@@
 Camera Height (m) $1.6\pm 0.05$ $1.0\pm 0.3$ $0.6\pm 0.8$ $0.9\pm 0.1$
-Focal Length (mm) $1146.8\pm 2.0$ $1172.4\pm 121.3$ $1962.2\pm 1.5$ $600\pm 0$
+Focal Length (mm) $146.8\pm 2.0 11$$172.4\pm 121.3$ $1962.2\pm 1.5$ $600\pm 0$
 No. of Joints 38 34 24 24
@@
 “Lifting Networks”, are widely used for their flexibility, augmentation capabilities, and ease of implementation.
-Given that 2D keypoint-based representations lack many visual cues present in images (such as occlusion), it simplifies the task of 3D pose estimation–often at the cost of generalization.
-As reported by [35, 20], they show poor performance when tested against other datasets.
+Given that 2D keypoint-based representations lack many visual cues present in images (such as occlusion), it simplifies the task of 3D pose estimation–often at the cost of generalization [36, 21].
 Therefore, we survey the lifting networks reported by previous works [22, 43, 45, 40, 42, 2, 7, 9, 46, 44, 41, 5, 49] and pick 18 configurations based on their availability and adaptability.
@@
 We also gathered five transformer-based models [46, 44, 45, 26, 14], two graph convolution models [40, 43], a spatio-temporal encoder model [42], and a diffusion-based model [27].
-Given that various models from the authors work were missing checkpoints or were trained on 17 joint skeleton configurations, we decided to standardize testing by retraining these models from scratch on the more common 16 joint skeleton configuration to remain fair and consistent across comparisons.
+Given that some of these works omitted some checkpoints or were trained on 17 joint skeleton configurations, we decided to standardize testing by retraining these models from scratch on the more common 16 joint skeleton configuration to remain fair and consistent across comparisons.
 We further ensure the retrained model is as good as the original one by comparing the reported same-dataset results directly.
@@
 Additionally, most models standardize the pose by centering the skeleton at the hip joint.
-As shown in table 1 and discussed in [35], datasets differ significantly in terms of joint definitions, camera intrinsics/extrinsics, image resolution, and other parameters.
+As shown in Tab.˜1 and discussed in [36], datasets differ significantly in terms of joint definitions, camera intrinsics/extrinsics, image resolution, and other parameters.
 As such, each dataset requires tailored preprocessing pipelines.
@@
 The H36M [11] dataset provides four cameras with their intrinsic and extrinsic matrices.
-While the intrinsic matrices stay the same for all 15 action sequences across 11 subjects, the extrinsic parameters (e.g.
-orientation and translation) change.
+While the intrinsic matrices stay the same for all 15 action sequences across 11 subjects, the extrinsic parameters (e.g., orientation and translation) change.
 There are 15 actions, 2 subactions per action, and 4 cameras, all acted by 11 subjects.
@@
 3.4 Framework Initialization
-[…]
+Our framework consists of five key modules, as illustrated in Fig.˜4.
+Experiments are orchestrated through a global configuration file in YAML format, specifying critical parameters such as model_type, num_workers, trained_on_normalized_data, output_3d, video_mode, num_joints, and num_frames, among others.
+The framework supports model checkpoints saved in either JIT or ONNX format.
+The trained_on_normalized_data flag indicates whether the model was trained on normalized data, while the video_mode toggle enables inference on multiple frames simultaneously, which is essential for temporal models.
+By adopting this configuration-driven approach, models can be seamlessly loaded as abstract entities, simplifying the complexity associated with conducting diverse experiments.
+Specifically, the Model Module initializes the chosen model, and the Dataset Module processes and prepares the datasets according to the configuration settings.
+The selection of joints used by the Dataset Module is determined by the num_joints parameter provided in the configuration.
+To maintain consistency across experiments, we primarily employ a standardized 16-joint ordering: hip, right hip, right knee, right ankle, left hip, left knee, left ankle, spine, neck, head, left shoulder, left elbow, left wrist, right shoulder, right elbow, and right wrist.
+For certain experiments, we alternatively employ a 14-joint subset, excluding the head and spine.
+All evaluated models require either screen-space normalization or z-score standardization.
+Screen-space normalization scales pixel coordinates from the original image dimensions $(0,w)$ (width) and $(0,h)$ (height) to a unit interval $(0,1)$.
+In contrast, z-score standardization involves normalizing joint coordinates, $X$ with $\frac{X-\mu}{\sigma}$, where $\mu$ is the mean and $\sigma$ is the standard deviation.
 4 Experiments and Results
@@
 4.2 Cross-Dataset Evaluation
-We report the MPJPE scores in table 2 and PA-MPJPE scores in table 3. We also include two cross-dataset results available from prior work marked with $\ddagger$ in the tables.
+We report the MPJPE scores in Tab.˜2 and PA-MPJPE scores in Tab.˜3. We also include two cross-dataset results available from prior work marked with $\ddagger$ in the tables.
 […]
-Table 2: Cross-dataset evaluation sorted by decreasing average MPJPE (mm).
+Table 2: Cross-dataset evaluation sorted by decreasing average MPJPE ($m$).
 All models are trained on H36M.
@@
 Cross-Dataset Evaluation
-Model Name H36M GPA 3DPW SURREAL Average Error (MPJPE $\downarrow$)
+Model Name Human3.6M GPA 3DPW SURREAL Average Error (MPJPE $\downarrow$)
 GraFormer [45] $\bullet$ 36.44 259.11 308.96 150.46 188.74
@@
 MixSTE [42] $\bullet$ 38.44 182.13 171.28 131.23 130.77
+DDHPose [1] $\bullet$ 38.28 200.29 138.64 129.15 126.59
 D3DP [27] $\bullet$ 39.61 189.74 148.56 127.90 126.45
@@
 […]
-Table 3: Cross-dataset evaluation sorted by decreasing average PA-MPJPE (mm).
+Table 3: Cross-dataset evaluation sorted by decreasing average PA-MPJPE ($m$).
 All models are trained on H36M.
@@
 Cross-Dataset Evaluation
-Model Name H36M GPA 3DPW SURREAL Average Error (PA-MPJPE $\downarrow$)
+Model Name Human3.6M GPA 3DPW SURREAL Average Error (PA-MPJPE $\downarrow$)
 SEM-GCN [9] $\diamond$ 36.12 178.43 166.76 87.35 117.17
@@
 Martinez [9] $\diamond$ 31.80 124.17 111.30 67.97 83.81
+DDHPose [1] $\bullet$ 30.13 139.24 76.85 80.95 81.79
 VideoPose [9] $\diamond$ 32.17 126.61 102.65 65.67 81.78
@@
 Figure 5: Per-joint errors for Protocols 1 and 2, averaged across all models and datasets.
-We also analyze per-joint position errors under both protocols in fig.
-5, averaged across all 18 model configurations and all four datasets.
+We also analyze per-joint position errors under both protocols in Fig.˜5, averaged across all 18 model configurations and all four datasets.
 While the endpoints of the human body—such as the wrists, ankles, and head—are more susceptible to errors, note that Procrustes alignment alone results in a substantial reduction in these errors (highlighted in red).
@@
 In the context of pose estimation, where datasets may originate from diverse distributions, such normalization indeed improves cross-dataset performance when the mean and standard deviation are known—though they are typically unavailable during inference time in real-world scenarios.
-Table 4 shows the models [22] and [42] trained with H36M and normalized with the mean and standard deviation of the same dataset.
+Table˜4 shows the models [22] and [43] trained with H36M and normalized with the mean and standard deviation of the same dataset.
 For the other datasets, we used their respective mean and standard deviation to normalize them.
-We also include the unoptimized ($\diamond$) and PoseAug-optimized ($\dagger$) variants from table 2 for comparison, and also note the percentage improvement relative to the unoptimized variant in both cases.
+We also include the unoptimized ($\diamond$) and PoseAug-optimized ($\dagger$) variants from Tab.˜2 for comparison, and also note the percentage improvement relative to the unoptimized variant in both cases.
 From all the cases, note that Z-score normalization performed with the test set’s mean and standard deviation improves cross-dataset generalization significantly for both models (e.g., SemGCN sees an improvement of more than 50% in 3DPW).
@@
 This shows that models overfit on the distribution they are trained on—which can often be countered by increasing variation.
-Table 4: MPJPE comparison of Martinez and SEM-GCN models across datasets. Percentage improvements relative to the baseline ($\diamond$) are in parentheses.
+Table 4: MPJPE comparison of Martinez and SEM-GCN models across datasets.
+Percentage improvements relative to the baseline ($\diamond$) are in parentheses.
+Arrows in green indicate improvement over the Unoptimized (baseline), and red arrows indicate degradation from baseline.
 (a) Martinez Comparison
 Model H36M GPA 3DPW SURREAL Average Error
-Unoptimised $\diamond$ 41.42 205.62 226.20 110.01 145.81
-Optimised $\dagger$ 39.11 ($\downarrow$ 5.6%) 169.79 ($\downarrow$ 17.4%) 134.12 ($\downarrow$ 40.7%) 98.99 ($\downarrow$ 10.0%) 110.50 ($\downarrow$ 24.2%)
-Z-score 52.37 ($\uparrow$ 26.4%) 104.39 ($\downarrow$ 49.3%) 141.10 ($\downarrow$ 37.6%) 81.64 ($\downarrow$ 25.8%) 94.88 ($\downarrow$ 34.9%)
+Unoptimized $\diamond$ 41.42 205.62 226.20 110.01 145.81
+Optimized $\dagger$ 39.11 ($\downarrow$ 5.6%) 169.79 ($\downarrow$ 17.4%) 134.12 ($\downarrow$ 40.7%) 98.99 ($\downarrow$ 10.0%) 110.50 ($\downarrow$ 24.2%)
+Z-score Normalization 52.37 ($\uparrow$ 26.4%) 104.39 ($\downarrow$ 49.3%) 141.10 ($\downarrow$ 37.6%) 81.64 ($\downarrow$ 25.8%) 94.88 ($\downarrow$ 34.9%)
 (b) SEM-GCN Comparison
 Model H36M GPA 3DPW SURREAL Average Error
-Unoptimised $\diamond$ 47.03 262.34 315.31 118.30 185.75
-Optimised $\dagger$ 41.90 ($\downarrow$ 10.9%) 241.21 ($\downarrow$ 8.1%) 239.07 ($\downarrow$ 24.2%) 107.26 ($\downarrow$ 9.3%) 157.36 ($\downarrow$ 15.3%)
-Z-score 53.94 ($\uparrow$ 14.7%) 114.85 ($\downarrow$ 56.2%) 153.61 ($\downarrow$ 51.3%) 99.88 ($\downarrow$ 15.6%) 105.57 ($\downarrow$ 43.2%)
+Unoptimized $\diamond$ 47.03 262.34 315.31 118.30 185.75
+Optimized $\dagger$ 41.90 ($\downarrow$ 10.9%) 241.21 ($\downarrow$ 8.1%) 239.07 ($\downarrow$ 24.2%) 107.26 ($\downarrow$ 9.3%) 157.36 ($\downarrow$ 15.3%)
+Z-score Normalization 53.94 ($\uparrow$ 14.7%) 114.85 ($\downarrow$ 56.2%) 153.61 ($\downarrow$ 51.3%) 99.88 ($\downarrow$ 15.6%) 105.57 ($\downarrow$ 43.2%)
 4.4 Impact of Viewpoint
 One interesting experiment is to look at the interplay between the viewpoint distribution of the datasets and its impact on the MPJPE error.
-In figs.
-6 and 7, we present the MPJPE scores against the viewpoint distribution over the training and testing sets simultaneously.
+In Figs.˜6 and 7, we present the MPJPE scores against the viewpoint distribution over the training and testing sets simultaneously.
 One common trend across all these figures is that whenever the viewpoint distribution disagrees with the training set’s viewpoint distribution, we observe a spike in error.
 This shows how much critical viewpoint is in 3D human pose estimation.
-Table 5 formally measures the inverse relationship between training data frequency and joint error by computing the Spearman correlation between the two, measured across patches on the viewpoint sphere measuring 5°5  ° $\times$ 10°10°10  ° degrees in elevation and azimuth, respectively.
+Table˜5 formally measures the inverse relationship between training data frequency and joint error by computing the Spearman correlation between the two, measured across patches on the viewpoint sphere measuring 5° $\times$ 10° degrees in elevation and azimuth, respectively.
 The correlation is always negative with high statistical significance.
 Notice how the p-value is lower when the evaluation is cross-dataset, indicating the variance in viewpoint distribution severely impacts performance.
-Please refer to Section A.2 for more figures.
+Please refer to Figure˜14 and Figure˜15 for greater comparison of error rates across models and datasets.
 (a) H36M
@@
 Figure 7: MPJPE score (SEM-GCN [43] trained on H36M) compared against viewpoint distribution in elevation relative to the subject.
-Table 5: Spearman correlation (which contained at least 5 training images and at least 5 test images) between training viewpoint distributions and test error (5°5  ° elev × 10°10°10  ° azim bins).
+Table 5: Spearman correlation (which contained at least 5 training images and at least 5 test images) between training viewpoint distributions and test error (5° elev × 10° azim bins).
 Num = number of bins; sigma is the p-value represented as the number of standard deviations from random.
 Num Train Test Spearman P-val Sigma
-[…]
-[…]
-[…]
-[…]
-[…]
-380 SURREAL H36M -0.19 0.018 380.018380.018\,380.018 38 3.70
-[…]
-[…]
-[…]
-[…]
-[…]
-417 H36M 3DPW -0.20 0.003 7010.0037010.003\,7010.003 701 4.13
-[…]
-[…]
-[…]
-[…]
+377 3DPW GPA -0.45 1.1e181.110-18 9.78
+380 3DPW H36M -0.30 2.4e072.410-07 6.10
+388 3DPW SURREAL -0.64 1.1e471.110-47 16.38
+304 3DPW 3DPW -0.44 2.8e142.810-14 8.55
+377 SURREAL GPA -0.33 2.6e092.610-09 6.85
+380 SURREAL H36M -0.19 0.018 380.018\,38 3.70
+388 SURREAL SURREAL -0.47 1.2e211.210-21 10.55
+304 SURREAL 3DPW -0.35 1.5e081.510-08 6.58
+641 H36M GPA -0.36 3.4e193.410-19 9.78
+939 H36M H36M -0.55 1.9e781.910-78 20.36
+891 H36M SURREAL -0.61 7.5e967.510-96 22.90
+417 H36M 3DPW -0.20 0.003 7010.003\,701 4.13
+621 GPA GPA -0.48 1.3e361.310-36 13.71
+738 GPA H36M -0.63 3.0e873.010-87 22.02
+751 GPA SURREAL -0.68 6.8e1126.810-112 25.50
+416 GPA 3DPW -0.39 1.3e141.310-14 8.58
+4.5 Correlation between Viewpoint and Error
+In Figs.˜6 and 7, we presented the error distribution curve with respect to the elevation and azimuth distribution separately.
+However, azimuth and elevation alone do not dictate the error—the number of training figures at a given point on the sphere dictates the error (i.e., both elevation and azimuth together).
+Therefore, we include additional contour plots showing the viewpoint and error distribution.
+Table˜5 showed a strong inverse correlation between the error and viewpoint distribution with high statistical significance.
+This is not observed in individual azimuth plots (e.g., Fig.˜6).
+In the contour plots (e.g., Figs.˜9(b), 11(b) and 13(b)), the viewpoint and error distribution are marked with green and red heatmaps.
+The X and Y axes represent azimuth and elevation, respectively.
+Whenever there is a trough in the viewpoint distribution—i.e.
+the training samples are under-sampled—the error goes up.
+This shows how much critical viewpoint is in 3D human pose estimation.
+Figure 9: (a) Viewpoint distribution vs.
+MPJPE ($m$) error.
+(b) Contour plot showing the error (in red) and viewpoint distribution (in green) with the x and y axes as azimuth and elevation, respectively.
+For both figures, errors are obtained on the GPA test set from [43] trained with the H36M dataset.
+Figure 11: (a) Viewpoint distribution vs.
+MPJPE ($m$) error.
+(b) Contour plot showing the error (in red) and viewpoint distribution (in green) with the x and y axes as azimuth and elevation, respectively.
+For both figures, errors are obtained on the 3DPW test set from [43] trained with the H36M dataset.
+Figure 13: (a) Viewpoint distribution vs.
+MPJPE ($m$) error.
+(b) Contour plot showing the error (in red) and viewpoint distribution (in green) with the x and y axes as azimuth and elevation, respectively.
+For both figures, errors are obtained on the SURREAL test set from [43] trained with the H36M dataset.
+Figure 14: Protocol 1 visual results from Tab.˜2 shown in increasing average dataset error.
+Figure 15: Protocol 2 visual results from Tab.˜3 shown in increasing average dataset error.
 5 Conclusion and Future Work
-We introduced a framework for evaluating cross-dataset performance of 2D-to-3D pose lifting networks.
+We have introduced a framework for evaluating cross-dataset performance of 2D-to-3D pose lifting networks.
 PoseBench3D enables standardized, reproducible evaluation and supports automatic benchmarking across multiple datasets.
@@
 Ultimately, we hope PoseBench3D will serve as a foundation for fair, comprehensive, and scalable benchmarking in 3D human pose estimation.
-Appendix A Experiments
-A.1 Example User-Supplied Config File
-Listing 1: User-Supplied YAML Config File
-# Unified YAML configuration
-# General settings
-dataset: [h36m, surreal, 3dpw, gpa]
-keypoints: ["gt": ground truth]
-checkpoint: "checkpoint"
-evaluate: "/path_to_checkpoint.pt"
-model_type: [JIT-Traced PyTorch File, ONNX TensorFlow File]
-# Dataset locations
-path_to_h36m: path_to_h36m
-path_to_gpa: path_to_gpa
-path_to_surreal_train: path_to_surreal_train
-path_to_surreal_test: path_to_surreal_test
-path_to_3dpw: path_to_3dpw
-data_dir: path_to_data_directory
-# Runtime & miscellaneous options
-num_workers: 2-4
-print_sample: [True, False]
-per_joint_error: [True, False]
-save_predictions: [True, False]
-# Input tensor shape
-input_shape:
-batch_size: [1…N]
-num_frames: 1
-num_joints: [14, 16]
-coordinates: 2
-# Output tensor shape
-output_shape:
-batch_size: [1…N]
-num_frames: 1
-num_joints: [14, 16]
-coordinates: 3
-# Model-specific flags
-model_info:
-flattened_coordinate_model: [True, False]
-output_3d: [Meters, Millimeters]
-root_center_2d_test_input: [True, False]
-normalize_2d_to_minus_one_to_one: [True, False]
-trained_on_normalized_data: [True, False]
-video_model: [True, False]
-A.2 Cross-Dataset Evaluation Visualization
-Figure 8: Protocol 1 visual results from table 2 shown in increasing average dataset error.
-Figure 9: Protocol 2 visual results from table 3 shown in increasing average dataset error.
-Figure 10: Per-joint errors for Protocols 1 and 2, averaged across all models on H36M.
-Figure 11: Per-joint errors for Protocols 1 and 2, averaged across all models on GPA.
-Figure 12: Per-joint errors for Protocols 1 and 2, averaged across all models on 3DPW.
-Figure 13: Per-joint errors for Protocols 1 and 2, averaged across all models on SURREAL.
-(a) H36M
-(b) GPA
-(c) 3DPW
-(d) SURREAL
-Figure 14: MPJPE score (SEM-GCN [42] trained on GPA) compared against viewpoint distribution in azimuth relative to the subject.
-(a) H36M
-(b) GPA
-(c) 3DPW
-(d) SURREAL
-Figure 15: MPJPE score (SEM-GCN [42] trained on GPA) compared against viewpoint distribution in elevation relative to the subject.
-(a) H36M
-(b) GPA
-(c) 3DPW
-(d) SURREAL
-Figure 16: MPJPE score (SEM-GCN [42] trained on 3DPW) compared against viewpoint distribution in azimuth relative to the subject.
-(a) H36M
-(b) GPA
-(c) 3DPW
-(d) SURREAL
-Figure 17: MPJPE score (SEM-GCN [42] trained on 3DPW) compared against viewpoint distribution in elevation relative to the subject.
-(a) H36M
-(b) GPA
-(c) 3DPW
-(d) SURREAL
-Figure 18: MPJPE score (SEM-GCN [42] trained on GPA) compared against viewpoint distribution in elevation relative to the subject.
-(a) H36M
-(b) GPA
-(c) 3DPW
-(d) SURREAL
-Figure 19: MPJPE score (SEM-GCN [42] trained on SURREAL) compared against viewpoint distribution in azimuth relative to the subject.
-(a) H36M
-(b) GPA
-(c) 3DPW
-(d) SURREAL
-Figure 20: MPJPE score (SEM-GCN [42] trained on SURREAL) compared against viewpoint distribution in elevation relative to the subject.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4.
-Justification: Both the abstract and introduction accurately reflect the paper’s scope.
-We discuss the current friction and research gap, how we set out to address these issues, introduce a benchmarking evaluation environment to partner our approach, and provide results according to our proposed approach.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: We discuss the limitation and future work in the conclusion section (section 5).
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: The paper does not contain theoretical results; instead, it focuses on the practical applications of human pose estimation.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19.
-Justification: The paper outlines in detail the requirements necessary to reproduce our results.
-We mention all three components: an extensive overview of required datasets section 3.1, the models and architectures supported by the framework section 3.2, and the user-supplied configuration file for customization section A.1.
-We make our codebase available to the public for extension and reproduction of our work.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24.
-Justification: Our framework is fully open-access, with a link to its repository provided.
-The repository includes detailed instructions for dataset setup, model import, and configuration file requirements needed to reproduce and extend our results.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29.
-Justification: We provide details about our experimental settings/details in the github repository linked with the paper.
-While we mention the specifics of our testing environment, our framework does not support training.
-For model training details and settings, please refer to the original work by the authors.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34. Justification: We present the statistical significance of the evaluation results against viewpoint distribution through Spearman correlation (c.f. table 5).
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: We present the compute resources in section 4.1.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44.
-Justification: Our paper clearly details the code and resources used in this project to ensure both the reproducibility of our results and ease of extension for future work.
-All raw datasets are publicly available at the original sources provided by their respective authors.
-Additionally, we provide the post-processed data used in our experiments.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49.
-Justification: We highlight the positive impact of advancing 3D Human Pose Estimation, particularly in rapidly growing fields as noted in the Introduction.
-We do not believe there are any significant negative societal impacts associated with improving 3D Human Pose Estimation.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: We expect that 3D Human Pose Estimation benchmarking and the datasets have a low risk for misuse.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59.
-Justification: We use four datasets, all of which are free for research purposes and credited with appropriate citations.
-The eighteen model configurations that we use for experiments or re-evaluation are licensed under one of the following: CC-BY-4.0, MIT, Apache License 2.0.
-Some of the models do not mention any citation.
-Nevertheless, we cited all the models used in our experiments.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: We provide the original code repository, checkpoints, and data preparation script with proper documentation.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: Our work does not make use of crowdsourcing or human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: Our work does not make use of human subjects
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [N/A]
-  79. Justification: Our work does not make use of LLMs in a way that is important to the core methods of this paper.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.